the sum of two number is 13. two times a second number minus three times the second number is 1. if you let x stand for the first number and y for the second number, what are the two numbers?
Answer by ankor@dixie-net.com(22740)

The problem seems to have an error, I assume you meant:
"two times a second number minus three times the first number is 1."
:
Write an equation for each statement.
:
"the sum of two numbers is 13."
x + y = 13
or
x = (13 - y)
or
y = (13 - x)
:
"two times a second number minus three times the first number is 1"
2y - 3x = 1
:
what are the two numbers?
:
Substitute (13-x) for y in the above equation and you have:
2(13-x) - 3x = 1
26 - 2x - 3x = 1
-5x = 1 - 26
x = -25/-5
x = + 5
:
y = (13 - 5) = 8
:
:
Check solution using the statement:
"two times a second number minus three times the first number is 1. "
2(8) - 3(15) = 1
:
If this is not what you meant, let me know, the problem can be done that way
You would have:
2y - 3y = 1
-y = 1
y = -1
:
Then:
x + (-1) = 13
x = 13 + 1
x = 14